### v1

# First variation of flat traces on negatively curved surfaces

Abstract.
For a closed negatively curved surface $(X,g)$ the flat trace of the geodesic Koopman operators $V_{g}^{\tau}f=f\circ G_{g}^{\tau}$ is the periodic orbit distribution
$\mathrm{Tr}^{\flat}V_{g}(\tau)=\sum_{\gamma}\frac{L_{\gamma}^{\#}}{\lvert\det(I-P_{\gamma})\rvert}\,\delta(\tau-L_{\gamma}),\qquad\tau>0,$
supported on the length spectrum and weighted by the linearized Poincaré maps $P_{\gamma}$. For a smooth family of negatively curved metrics $g_{t}$ we compute the first variation $\partial_{t}|_{0}\,\mathrm{Tr}^{\flat}V_{g_{t}}$ as a distribution. At an isolated length $\ell$ the leading singularity is a multiple of $\delta^{\prime}(\tau-\ell)$, and its coefficient is an explicit linear functional of the length variations $\dot{L}_{\gamma^{m}}$ of the closed geodesics with $L_{\gamma^{m}}=\ell$. This transport coefficient forces the marked lengths to be locally constant along any deformation with constant flat trace. As an application, if $\mathrm{Tr}^{\flat}V_{g_{t}}=\mathrm{Tr}^{\flat}V_{g_{0}}$ for all $t$ then $g_{t}$ is isometric to $g_{0}$ for all $t$. Together with Sunada-type constructions of non isometric pairs with equal flat traces, this shows that the flat trace is globally non-unique yet locally complete along smooth families.
Key words and phrases:
flat trace, dynamical zeta, negative curvature, Anosov geodesic flow, Livšic theorem
2020 Mathematics Subject Classification:
37D40, 58J50, 37C30, 53C24
Research partially supported by NSF RTG grant DMS-1502632.
1. Introduction
The dynamical side of spectral geometry rests on periodic orbit expansions attached to the geodesic flow. On a closed negatively curved surface, the flat trace of the geodesic Koopman operators $V_{g}^{\tau}f=f\circ G_{g}^{\tau}$ is the distribution supported on the length spectrum
(1) $\mathrm{Tr}^{\flat}V_{g}(\tau)=\sum_{\gamma}\frac{L_{\gamma}^{\#}}{\lvert\det(I-P_{\gamma})\rvert}\,\delta(\tau-L_{\gamma}).$
Its Laplace transform gives the logarithmic derivative of the Guillemin Ruelle dynamical zeta built with the same weights, so equality of flat traces implies equality of the corresponding zetas.
Let $X$ be a closed oriented surface and let $g_{t}$ be a $C^{\infty}$ family of negatively curved metrics on $X$ for $t\in(-\epsilon,\epsilon)$. Write $G_{t}^{\tau}$ for the geodesic flow on $S^{*}_{g_{t}}X$, write $V_{t}^{\tau}f=f\circ G_{t}^{\tau}$ for the Koopman operator on $L^{2}(S^{*}_{g_{t}}X,d\mu_{L,t})$, and let $\mathrm{Tr}^{\flat}V_{t}$ denote its flat trace distribution on $(0,\infty)$. Here $d\mu_{L,t}$ is the Liouville measure induced on $S^{*}_{g_{t}}X$ by the Hamiltonian $H_{t}(x,\xi)=\frac{1}{2}\lvert\xi_{x}\rvert_{g_{t}}^{2}$.
The main analytic input is the distributional first variation of $\mathrm{Tr}^{\flat}V_{t}$ under a metric deformation. At an isolated length $\ell>0$, the distribution $\partial_{t}|_{0}\,\mathrm{Tr}^{\flat}V_{t}$ has a leading singularity of the form $\mathcal{T}(\ell)\,\delta^{\prime}(\tau-\ell)$. The coefficient $\mathcal{T}(\ell)$ is an explicit transport functional of the marked length variations. Proposition 2.2 gives
(2) $\mathcal{T}(\ell)=-\sum_{\begin{subarray}{c}\gamma,m\\ L_{\gamma^{m}}=\ell\end{subarray}}\frac{L_{\gamma}^{\#}}{\lvert\det(I-P_{\gamma}^{m})\rvert}\,\dot{L}_{\gamma^{m}}.$
In particular, constancy of the flat trace forces $\mathcal{T}(\ell)=0$ for every isolated $\ell$.
Once the transport coefficients vanish, the persistence of hyperbolic periodic orbits under smooth perturbation implies $\dot{L}_{\gamma}=0$ for every closed geodesic. Equivalently the marked length spectrum is locally constant in $t$. At this point one may invoke the marked length spectrum rigidity theorem of Croke [1] and Otal [8] to conclude that $g_{t}$ is isometric to $g_{0}$ for all $t$. We continue instead with a direct deformation theoretic argument in the same dynamical framework. After the $\delta^{\prime}$ constraint fixes all periods, the Guillemin-Kazhdan variational identity yields vanishing of the period integrals of $\dot{p}=\partial_{t}|_{0}\,\lvert\xi\rvert_{g_{t}}$. Livšic theory converts these vanishing period integrals into a cohomological equation on $S^{*}_{g_{0}}X$, and the Guillemin-Kazhdan operator calculus identifies $\dot{p}$ with a Lie derivative on the base. This completes a proof of infinitesimal triviality which does not appeal to marked length spectrum rigidity and which is expected to be robust for Anosov flows in higher dimension.
Remark 1.1.
In dimension $n\geq 3$, the Lefschetz fixed point reduction and the Abel-type extraction of the $\delta^{\prime}$-term remain formally identical, but the clean codimension and the harmonic bookkeeping change. Concretely, the circle decomposition in Appendix 6.2 is replaced by the SO(n1)SO(n-1) representation theory of spherical harmonics on $\mathbb{S}^{n-2}$, so the mode reduction and coercive estimates must be organized by irreducible SO(n1)SO(n-1) types rather than Fourier modes of $SO(2)$. In this regime, the marked length spectrum rigidity input of Croke and Otal is not available, so the direct flat trace variational approach is the natural substitute once the SO(n1)SO(n-1)-mode calculus is implemented.
As an application of this analysis we obtain the following pathwise rigidity statement.
Theorem 1.2.
If $\mathrm{Tr}^{\flat}V_{g_{t}}=\mathrm{Tr}^{\flat}V_{g_{0}}$ as distributions on $(0,\infty)$ for all $t$, then there exists a smooth one parameter family of diffeomorphisms $\{\varphi_{t}\}$ of $X$ with $\varphi_{0}=\mathrm{id}$ such that
$g_{t}=\varphi_{t}^{*}g_{0}\qquad\text{for all }t\in(-\epsilon,\epsilon).$
Equivalently, $(X,g_{t})$ is isometric to $(X,g_{0})$ for every $t$.
Sunada type constructions produce non isometric pairs with equal flat traces at fixed metrics [6]. Together with those examples, Theorem 1.2 shows that the flat trace is globally non unique yet locally complete along smooth paths inside the space of negatively curved metrics.
Remark 1.3.
The rigidity statement in Theorem 1.2 is weaker than marked length spectrum rigidity, since it assumes a smooth deformation. The point of the present paper is the distributional first variation formula for the flat trace, in particular the transport coefficient in Proposition 2.2. This coefficient does not seem to appear in the literature in this explicit form, and it is the mechanism that converts constancy of a zeta type invariant into infinitesimal constraints on the marked lengths. The appendices collect auxiliary calculations and are included for completeness.
The paper is organized as follows. Section 2 recalls the flat trace distribution of the geodesic Koopman operator and derives its first variation. Section 3 records the metric variation of closed geodesic lengths and of the Hamiltonian $p_{t}(x,\xi)=\lvert\xi\rvert_{g_{t}}$. Section 4 expresses the $\delta^{\prime}$-coefficient explicitly in terms of the metric perturbation $h=\dot{g}$ and extracts the resulting linear relation among the $h$-period integrals. Section 5 combines the $\delta^{\prime}$ constraint with Livšic theory and the Guillemin-Kazhdan $SO(2)$ calculus, summarized in Appendix 6.2, to conclude infinitesimal triviality and then integrate it in $t$ to obtain Theorem 1.2. Appendix 6.1 contains the verification that the flat trace formula has no additional lower order singular terms beyond the Lefschetz coefficients.
Acknowledgements
This article refines and unifies material that appeared in the author’s doctoral dissertation at Northwestern University. The author gratefully acknowledges the guidance of the late Steve Zelditch.
2. Flat trace, clean fixed sets, and first variation
2.1. Kernel and flat trace as a pullback and pushforward
For $\tau>0$, the Schwartz kernel of $V^{\tau}_{g}$ is the $\delta$-section
$K^{\tau}_{t}(\zeta,\zeta^{\prime})=\delta(\zeta^{\prime}-G^{\tau}_{t}\zeta),\qquad\zeta,\zeta^{\prime}\in S^{*}_{g_{t}}X,$
understood as a conormal distribution to $\mathrm{Graph}(G^{\tau}_{t})$. The flat trace is the pullback by the diagonal $\iota:\zeta\mapsto(\zeta,\zeta)$ followed by fiber integration
(3) $\mathrm{Tr}^{\flat}V_{t}(\tau)=\int_{S^{*}_{g_{t}}X}\iota^{*}K^{\tau}_{t}(\zeta)d\mu_{L,t}(\zeta)=\int_{S^{*}_{g_{t}}X}\delta(\zeta-G^{\tau}_{t}\zeta)d\mu_{L,t}(\zeta).$
In negative curvature, the fixed set of $G^{\tau}_{t}$ is clean precisely when $\tau$ equals the period of a (possibly iterated) closed geodesic.
2.2. Normal form near a closed orbit
Fix a primitive closed geodesic $\gamma$ of $(X,g_{0})$ with prime length $L^{\#}_{\gamma}>0$, linearized Poincaré map $P_{\gamma}$ and let $\gamma^{m}$ denote its $m$-iterate of length $L_{\gamma^{m}}=mL^{\#}_{\gamma}$. For $\tau$ near $L_{\gamma^{m}}$, the fixed set of $G^{\tau}_{0}$ is the circle $\gamma$ (as a subset of $S^{*}_{g_{0}}X$). Choose symplectic coordinates $(s,y;\sigma,\eta)$ near $\gamma$ so that $s\in\mathbb{R}\backslash L^{\#}_{\gamma}\mathbb{Z}$ parametrizes the orbit, $(y,\eta)\in\mathbb{R}^{2}$ are transverse canonical coordinates, and the flow has the normal form
$G^{\tau}_{0}:(s,y;\sigma,\eta)\mapsto(s+\tau,\Phi_{\tau}(y,\eta);\sigma,\Psi_{\tau}(y,\eta)),$
$(y,\eta)=(0,0)$ on $\gamma$, and $d\Phi_{L_{\gamma^{m}}}$ conjugates to $P^{m}_{\gamma}$ on the transversal.
As a conormal distribution, $K^{\tau}_{0}$ may be represented by an oscillatory integral
$K^{\tau}_{0}(\zeta,\zeta^{\prime})=(2\pi)^{-N}\int e^{i\langle\theta,\zeta^{\prime}-G^{\tau}_{0}\zeta\rangle}a_{0}(\zeta,\zeta^{\prime},\tau,\theta)d\theta$
with a classical amplitude $a_{0}$. Pulling back by the diagonal and integrating over $\zeta$ (cf. (8)) gives
(4) $\mathrm{Tr}^{\flat}V_{0}(\tau)=(2\pi)^{-N}\int e^{i\langle\theta,\zeta-G^{\tau}_{0}\zeta\rangle}a_{0}(\zeta,\zeta,\tau,\sigma)d\sigma d\zeta.$
2.3. Stationary phase with a clean one dimensional critical set
The phase $\Phi_{\tau}(\zeta,\theta):=\langle\theta,\zeta-G^{\tau}_{0}\zeta\rangle$ has critical set
$\mathcal{C}_{\tau}=\{(\zeta,\theta):\zeta\in\mathrm{Fix}(G^{\tau}_{0}),\theta\in N^{*}_{\zeta}\mathrm{Fix}(G^{\tau}_{0})\},$
which, for $\tau$ near $L_{\gamma^{m}}$, is a vector bundle over the circle $\gamma$. The clean stationary-phase theorem ([2]) yields
(5) $\mathrm{Tr}^{\flat}V_{0}(\tau)\sim\sum_{m\geq 1}\sum_{\gamma}\int_{\mathbb{R}}e^{i\sigma(\tau-L_{\gamma^{m}})}b_{\gamma,m}(\sigma)d\sigma\quad+\quad C^{\infty}(\tau),$
where, crucially, the leading term of the classical symbol $b_{\gamma,m}$ is
(6) $b^{(0)}_{\gamma,m}=(2\pi)^{-1}\frac{L^{\#}_{\gamma}}{|\det(I-P^{m}_{\gamma})|}.$
By evaluating the $\sigma$-integral, we have
(7) $\mathrm{Tr}^{\flat}V_{t}(\tau)=\sum_{m\geq 1}\sum_{\gamma}\frac{L^{\#}_{\gamma}}{|\det(I-P^{m}_{\gamma})|}\delta(\tau-L_{\gamma^{m}})+\,(\text{lower order terms at $\tau=L_{\gamma^{m}}$})+C^{\infty}(\tau),$
which is the standard flat-trace formula whereby the subprincipal corrections in (5) yields only derivatives of $\delta$ of order $\leq 0$, i.e. multiples of $\delta$ because the critical set is 1-dimensional.
Remark 2.1.
For a Fourier integral operator on an $n$-dimensional manifold whose fixed set at time $\tau$ is clean of dimension $d$, the clean stationary phase theorem shows that the flat trace singularity at $\tau$ has order (nd1)/2(n-d-1)/2. In our setting, $n=\dim(S^{*}_{g}X)=3$ and the relevant fixed sets are the closed geodesics, so $d=1$. Hence, at a fixed metric $g_{t}$, the singularity at $\tau=L_{\gamma^{m}}(t)$ is of order 0, and no derivatives $\delta^{(k)}$ with $k\geq 1$ appear in the flat trace at fixed $t$. By contrast, differentiating with respect to the parameter $t$ may produce $\delta^{\prime}$-terms through transport of the singular support (cf. Proposition 2.2).
The absence of any further singular contributions at a fixed period (i.e. the fact that (7) has no additional “lower order” $\delta$-terms) is proved in Appendix 6.1.
2.4. Parameter dependent phase and $t$ variation of the clean contribution
Let $g_{t}$ be a small deformation. Let $\chi\in C^{\infty}_{c}$ localize near a single pair $(\gamma,m)$. Microlocally near $\gamma$,
(8) $\mathrm{Tr}^{\flat}V_{t}(\tau)=\int\int e^{i\Phi_{t}(\zeta,\theta,\tau)}a_{t}(\zeta,\theta,\tau)d\theta d\zeta,\quad\text{with}~~\Phi_{t}=\langle\theta,\zeta-G^{\tau}_{t}\zeta\rangle.$
As before, we apply the method of joint-normal stationary phase to the clean critical submanifold
(9) $C_{t,\tau=L_{\gamma^{m}}(t)}=\{(\zeta,\theta):\ \zeta\in\gamma_{t},\ \theta\in N_{\zeta}^{*}\gamma_{t}\}.$
to obtain
(10) $\mathrm{Tr}^{\flat}V_{t}(\tau)\sim\int_{\mathbb{R}}e^{i\sigma(\tau-L_{\gamma^{m}}(t))}\,b_{\gamma,m}(t,\sigma)d\sigma\quad+\quad C^{\infty}(\tau,t),$
with (cf. (6))
$b_{\gamma,m}(t,\sigma)\sim\sum_{j\geq 0}b^{(j)}_{\gamma,m}(t)\langle\sigma\rangle^{-j},\quad b^{(0)}_{\gamma,m}(t)=(2\pi)^{-1}\frac{L^{\#}_{\gamma}(t)}{|\det(I-P^{m}_{\gamma,t})|}$
For a test function $\psi\in C^{\infty}_{c}((0,\infty))$, we pair (10) with $\psi$ after differentiating at $t=0$ to obtain
(11) $\langle\partial_{t}\big|_{0}\mathrm{Tr}^{\flat}V_{t},\,\psi\rangle=\int_{\mathbb{R}}e^{-i\sigma L_{\gamma^{m}}}\check{\psi}(\sigma)\left(\partial_{t}b_{\gamma,m}(t,\sigma)\big|_{t=0}-i\sigma b_{\gamma,m}(0,\sigma)(m\dot{L}^{\#}_{\gamma})\right)d\sigma$
where $\check{\psi}(\sigma)=\int_{\mathbb{R}_{>0}}e^{i\sigma\tau}\psi(\tau)d\tau$ and $m\dot{L}^{\#}_{\gamma}=\partial_{t}(L_{\gamma^{m}}(t))\big|_{t=0}$. By the asymptotic expansion
$b_{\gamma,m}(0,\sigma)=(2\pi)^{-1}\frac{L^{\#}_{\gamma}}{|\det(I-P^{m}_{\gamma})|}+O(\langle\sigma\rangle^{-1})$
and the Fourier identities
$\int e^{i\sigma(\tau-\ell)}d\sigma=2\pi\delta(\tau-\ell),\quad\int e^{i\sigma(\tau-\ell)}(i\sigma)d\sigma=2\pi\delta^{\prime}(\tau-\ell),$
summing over all $(\gamma,m)$ provides, in the sense of distribution in $\tau$,
(12) $\partial_{t}\big|_{0}\mathrm{Tr}^{\flat}V_{t}(\tau)=-\sum_{\gamma}\sum_{m\geq 1}\frac{L_{\gamma}^{\#}}{|\det(I-P^{m}_{\gamma})|}(m\dot{L}^{\#}_{\gamma})\delta^{\prime}(\tau-L_{\gamma^{m}})+\sum_{\gamma,m}\dot{A}_{\gamma,m}\delta(\tau-L_{\gamma^{m}})$
where $\dot{A}_{\gamma,m}$ are smooth amplitude-variation weights stemming from $\partial_{t}b_{\gamma,m}(t,\sigma)\big|_{t=0}$, in particular, from $\partial_{t}\det(I-P^{m}_{\gamma})$, the Liouville density, etc. The first term is the most singular because it is the only source of $\delta^{\prime}$.
2.5. Isolation of the most singular terms and the weighted identities
Let $\ell>0$ be such that the set $\{(\gamma,m)\,:\,L_{\gamma^{m}}=\ell\}$ is finite (this is true in negative curvature). Take $\psi\in C^{\infty}_{c}((0,\infty))$ supported in a small neighborhood of $\ell$ and normalized by $\psi(\ell)=0,\quad\psi^{\prime}(\ell)=1$. We again pair (12) with $\psi$. Since $\langle\delta(\tau-\ell),\psi\rangle=\psi(\ell)=0$ and $\langle\delta^{\prime}(\tau-\ell),\psi\rangle=-\psi^{\prime}(\ell)=-1$, we get from (12)
(13) $\big\langle\partial_{t}\big|_{0}\mathrm{Tr}^{\flat}V_{t},\,\psi\big\rangle=\sum_{\begin{subarray}{c}\gamma,m\geq 1\\ L_{\gamma^{m}}=\ell\end{subarray}}\frac{L^{\#}_{\gamma}}{|\det(I-P^{m}_{\gamma})|}m\dot{L}_{\gamma}^{\#}.$
Therefore,
Proposition 2.2.
At each isolated length $\ell>0$, the length-transport coefficient of $\delta^{\prime}(\tau-\ell)$, in $\partial_{t}\big|_{0}\mathrm{Tr}^{\flat}V_{t}(\tau)$ equals
$\mathcal{T}(\ell):=-\sum_{\begin{subarray}{c}\gamma,m\\ L_{\gamma^{m}}=\ell\end{subarray}}\frac{L^{\#}_{\gamma}}{|\det(I-P^{m}_{\gamma})|}\dot{L}_{\gamma^{m}}.$
In particular, if $\mathrm{Tr}^{\flat}V_{t}$ is constant in $t$, then for every $\ell$, $\mathcal{T}(\ell)=0$.
Remark 2.3.
$\mathcal{T}(\ell)$ captures that the $\delta^{\prime}$ arises from transport of the singular support as the lengths move as a consequence of the metric deformation. In fact, no other term in (12) contributes to $\delta^{\prime}$ as the amplitude variations always yield $\delta$ (or smoother) at $\tau=\ell$. Secondly, $\mathcal{T}(\ell)=0$ due to constancy is the exact $\delta^{\prime}$-constraint needed in $\S$ 4. Namely, after identifying $m\dot{L}_{\gamma}^{\#}=\dot{L}_{\gamma^{m}}=\frac{1}{2}\int_{\gamma^{m}}h(T,T)ds$, it becomes a linear relation among the integrals of $h$ over orbits of length $\ell$. Lastly, when the flat trace is constant for all $t$, the support of $\mathrm{Tr}^{\flat}V_{t}$ is independent of $t$. Indeed, by structural stability and discreteness of the length spectrum, this forces each $L_{[\alpha]}(t)$ to be locally constant, hence constant, which is what we use subsequently .
3. The variations $\dot{L}_{\gamma}$ and $\dot{p}$ in terms of $\dot{g}$
Throughout this section, we fix $t=0$ and write
(14) $g_{t}=g+th+O(t^{2}),\quad h\in\Gamma(S^{2}T^{*}X).$
3.1. Variation of length for a closed geodesic
Fix a free homotopy class $\alpha]$ of loops in $X$, and for each $t$ let $\gamma_{t}:S^{1}\to X$ be the unique closed $g_{t}$-geodesic in $\alpha]$. We parametrize $\gamma_{t}$ with constant $g_{t}$-speed on $[0,1]$, i.e.
$|\dot{\gamma}_{t}(u)|_{g_{t}}\equiv L_{[\alpha]}(t),\qquad L_{[\alpha]}(t)=\int_{0}^{1}|\dot{\gamma}_{t}(u)|_{g_{t}}\,du.$
Write $\gamma:=\gamma_{0}$, $L:=L_{[\alpha]}(0)$ and set
$T:=\frac{\dot{\gamma}}{|\dot{\gamma}|_{g}}\in TX\big|_{\gamma},\qquad ds:=|\dot{\gamma}|_{g}\,du=L\,du,$
so that $T$ is the $g$-unit tangent vector field and $s$ is the $g$-arclength parameter.
Let $V:=\partial_{t}\gamma_{t}|_{t=0}$ be the variational vector field along $\gamma$. Differentiating
$L_{[\alpha]}(t)=\int_{0}^{1}\sqrt{g_{t}(\dot{\gamma}_{t},\dot{\gamma}_{t})}\,du$
at $t=0$ gives
$\dot{L}_{[\alpha]}(0)$ $=\frac{1}{2}\int_{0}^{1}\frac{h(\dot{\gamma},\dot{\gamma})+2g(\nabla_{u}V,\dot{\gamma})}{|\dot{\gamma}|_{g}}\,du$
(15) $=\frac{1}{2}\int_{0}^{1}h(T,T)\,ds+\int_{0}^{1}g(\nabla_{u}V,T)\,du.$
The second term vanishes by an integration by parts. Indeed, since
$\frac{d}{du}g(V,T)=g(\nabla_{u}V,T)+g(V,\nabla_{u}T),$
and $\gamma$ is a $g$-geodesic, $\nabla_{T}T=0$, hence $\nabla_{u}T=(ds/du)\nabla_{T}T=0$. Therefore $g(\nabla_{u}V,T)=\frac{d}{du}g(V,T)$, and by periodicity on $S^{1}$ we obtain
$\int_{0}^{1}g(\nabla_{u}V,T)\,du=g(V,T)\Big|_{u=0}^{u=1}=0.$
Substituting into (15) yields the standard first variation formula
(16) $\dot{L}_{[\alpha]}(0)=\frac{1}{2}\int_{\gamma}h(T,T)\,ds.$
Equivalently, for any closed geodesic $\gamma$ (and its $m$-iterate $\gamma^{m}$),
(17) $\dot{L}_{\gamma^{m}}=\frac{1}{2}\int_{\gamma^{m}}h(T,T)\,ds.$
3.2. Variation of the Hamiltonian on $T^{*}X$
Let $H_{t}:T^{*}X\to\mathbb{R}$ be the kinetic energy Hamiltonian
$H_{t}(x,\xi):=\frac{1}{2}|\xi|^{2}_{g_{t}}=\frac{1}{2}\,g_{t}^{ab}(x)\,\xi_{a}\xi_{b},$
and write
$p_{t}(x,\xi):=|\xi|_{g_{t}}=\sqrt{2H_{t}(x,\xi)}.$
Differentiating $g_{t}^{ab}g_{t,bc}=\delta^{a}_{c}$ gives $\dot{g}^{ab}=-h^{ab}$, where $h^{ab}:=g^{ac}g^{bd}h_{cd}$. Therefore
$\dot{H}=\frac{1}{2}\,\dot{g}^{ab}\xi_{a}\xi_{b}=-\frac{1}{2}\,h^{ab}\xi_{a}\xi_{b}.$
Since $p_{t}=(2H_{t})^{1/2}$, we obtain
(18) $\dot{p}=\frac{\dot{H}}{p}=-\frac{1}{2p}\,h^{ab}\xi_{a}\xi_{b}.$
In particular, on the unit cosphere bundle $S^{*}_{g}X=\{p=1\}$,
(19) $\dot{p}\big|_{S^{*}_{g}X}=-\frac{1}{2}\,h^{ab}\xi_{a}\xi_{b}=-\frac{1}{2}\,h(T,T),$
where $T=\xi^{\sharp}\in S_{g}X$ is the corresponding unit tangent vector. Combining (19) with (17) gives, for every closed geodesic $\gamma$,
(20) $\dot{L}_{\gamma}=-\int_{\gamma}\dot{p}\,ds.$
Moreover, $\dot{p}$ is an even function under the fiber flip $\xi\mapsto-\xi$, and in the $SO(2)$-Fourier decomposition on $S_{g}X$ its Fourier support is contained in modes $0,\pm 2$ (cf. Appendix 6.2).
4. Variation of the flat trace with explicit coefficients
This section records the explicit form of the $\delta^{\prime}$-constraint in geometric terms. For each fixed $t$ the flat trace of the Koopman operator of the geodesic flow is given by the (exact) Lefschetz/Guillemin formula
(21) $\mathrm{Tr}^{\flat}V_{t}(\tau)=\sum_{\gamma\in\mathcal{P}(g_{t})}\ \sum_{m\geq 1}\frac{L_{\gamma}^{\#}(t)}{\left|\det\!\left(I-P_{\gamma,t}^{\,m}\right)\right|}\,\delta\!\big(\tau-L_{\gamma^{m}}(t)\big)\;+\;C^{\infty}(\tau),$
where $\mathcal{P}(g_{t})$ denotes the set of primitive closed $g_{t}$-geodesics, $L_{\gamma^{m}}(t)=mL^{\#}_{\gamma}(t)$ is the length of the $m$-iterate, and $P_{\gamma,t}$ is the linearized Poincaré map (cf. Appendix 6.1).
Differentiating (21) at $t=0$ gives a distribution supported on the length spectrum. The only possible $\delta^{\prime}$-terms come from differentiating the moving supports $\tau=L_{\gamma^{m}}(t)$, and one obtains (cf. Proposition 2.2)
(22) $\partial_{t}\Big|_{t=0}\mathrm{Tr}^{\flat}V_{t}(\tau)=-\sum_{\gamma,m}\frac{L^{\#}_{\gamma}}{\left|\det\!\left(I-P_{\gamma}^{\,m}\right)\right|}\,\dot{L}_{\gamma^{m}}\,\delta^{\prime}\!\big(\tau-L_{\gamma^{m}}\big)\;+\;\text{\rm(only $\delta$ and smoother)}.$
Combining (22) with the length variation formula (17) yields an explicit expression for the $\delta^{\prime}$-coefficient in terms of $h=\dot{g}$:
(23) $\text{Coeff}_{\delta^{\prime}(\tau-L_{\gamma^{m}})}\Big(\partial_{t}\mathrm{Tr}^{\flat}V_{t}\Big|_{t=0}\Big)=-\frac{L^{\#}_{\gamma}}{\left|\det\!\left(I-P_{\gamma}^{\,m}\right)\right|}\,\frac{1}{2}\int_{\gamma^{m}}h(T,T)\,ds.$
In particular, if $\mathrm{Tr}^{\flat}V_{t}$ is constant in $t$, then for every $\ell$ in the length spectrum,
(24) $\sum_{\begin{subarray}{c}\gamma,m\\ L_{\gamma^{m}}=\ell\end{subarray}}\frac{L^{\#}_{\gamma}}{\left|\det\!\left(I-P_{\gamma}^{\,m}\right)\right|}\,\int_{\gamma^{m}}h(T,T)\,ds=0.$
When the length spectrum is simple this already implies $\int_{\gamma}h(T,T)\,ds=0$ for each closed $\gamma$; in general we will use instead the stronger support argument in §5.2 to deduce that all lengths are constant in $t$.
5. Triviality from the vanishing of $\delta^{\prime}$
5.1. Guillemin-Kazhdan variational identity
Let $p_{t}$ be as in $\S$ 3.2 and $y_{t}\subset\{p_{t}=1\}$ be the closed characteristic corresponding to $\gamma_{t}$. If $L_{\gamma_{t}}$ is constant in $t$ (see $\S$ 5.2), then Guillemin-Kazhdan’s strip argument yields
(25) $\int_{\gamma}\dot{p}ds=0$
Lemma 5.1 (Guillemin-Kazhdan).
Suppose $t\mapsto y_{t}$ is a $C^{\infty}$ family of closed orbits of the Hamiltonian flows of $p_{t}$ on $\{p_{t}=1\}$, all with the same period $L$. Then
$\int_{y_{0}}\dot{p}\,ds=0,$
where $ds$ denotes the Hamiltonian time parameter along $y_{0}$.
Proof.
Let $\alpha$ be the canonical $1$-form on $T^{*}X$ and $\omega=d\alpha$ the canonical symplectic form. Let $H_{t}$ be the Hamiltonian vector field of $p_{t}$, characterized by
$\iota_{H_{t}}\omega=dp_{t}.$
Choose a parametrization $\Phi:[0,\varepsilon]\times(\mathbb{R}/L\mathbb{Z})\to T^{*}X$ of the cylinder swept out by the closed orbits,
$\Phi(t,s):=y_{t}(s),\qquad\partial_{s}\Phi(t,s)=H_{t}(\Phi(t,s)).$
Since $p_{t}\circ\Phi\equiv 1$, differentiating in $t$ gives
$0=\partial_{t}(p_{t}\circ\Phi)=\dot{p}_{t}(\Phi)+dp_{t}(\partial_{t}\Phi).$
Using $\omega(H_{t},\cdot)=dp_{t}(\cdot)$, we compute
$\omega(\partial_{t}\Phi,\partial_{s}\Phi)=\omega(\partial_{t}\Phi,H_{t})=-\,\omega(H_{t},\partial_{t}\Phi)=-\,dp_{t}(\partial_{t}\Phi)=\dot{p}_{t}(\Phi).$
Hence $\Phi^{*}\omega=\dot{p}_{t}(\Phi)\,dt\wedge ds$. By Stokes’ theorem,
$\int_{0}^{\varepsilon}\int_{0}^{L}\dot{p}_{t}(y_{t}(s))\,ds\,dt$ $=\int_{[0,\varepsilon]\times(\mathbb{R}/L\mathbb{Z})}\Phi^{*}\omega=\int_{[0,\varepsilon]\times(\mathbb{R}/L\mathbb{Z})}d(\Phi^{*}\alpha)$
$=\int_{\partial([0,\varepsilon]\times(\mathbb{R}/L\mathbb{Z}))}\Phi^{*}\alpha=\int_{y_{\varepsilon}}\alpha-\int_{y_{0}}\alpha.$
Since $p_{t}$ is homogeneous of degree $1$ in $\xi$, Euler’s identity gives $\alpha(H_{t})=p_{t}$, and therefore
$\int_{y_{t}}\alpha=\int_{0}^{L}\alpha(H_{t})\,ds=\int_{0}^{L}p_{t}\,ds=L,$
which is independent of $t$ by hypothesis. Thus the right-hand side vanishes, and differentiating the left-hand side at $\varepsilon=0$ yields $\int_{y_{0}}\dot{p}\,ds=0$. ∎
5.2. Deducing constant lengths from the $\delta^{\prime}$ vanishing
Assume that $\mathrm{Tr}^{\flat}V_{g_{t}}=\mathrm{Tr}^{\flat}V_{g_{0}}$ as distributions on $(0,\infty)$ for all $t\in(-\varepsilon,\varepsilon)$. By the exact flat trace formula (21), for each fixed $t$ we have
$\mathrm{Tr}^{\flat}V_{g_{t}}(\tau)=\sum_{\gamma\in\mathcal{P}(g_{t})}\ \sum_{m\geq 1}\frac{L^{\#}_{\gamma}(t)}{\left|\det\!\left(I-P_{\gamma,t}^{\,m}\right)\right|}\,\delta\!\big(\tau-L_{\gamma^{m}}(t)\big)\;+\;C^{\infty}(\tau).$
Every coefficient in this expansion is strictly positive, hence the singular support is exactly the set of periods of closed orbits (the length spectrum with iterates):
$\mathrm{sing\,supp}\big(\mathrm{Tr}^{\flat}V_{g_{t}}\big)=\{L_{\gamma^{m}}(t):\gamma\in\mathcal{P}(g_{t}),\ m\geq 1\}.$
Since $\mathrm{Tr}^{\flat}V_{g_{t}}$ is independent of $t$, its singular support is independent of $t$ as a subset of $(0,\infty)$. Therefore the length spectrum of $g_{t}$ (including iterates) is locally constant as a set.
Fix a free homotopy class $\alpha]$. For each $t$ there is a unique closed $g_{t}$-geodesic $\gamma_{[\alpha]}(t)$ in this class, and by structural stability of Anosov flows the corresponding periodic orbit persists and varies continuously with $t$. In particular, the length function $t\mapsto L_{[\alpha]}(t)$ is continuous. But by the previous paragraph, $L_{[\alpha]}(t)$ takes values in the fixed set $\mathrm{sing\,supp}(\mathrm{Tr}^{\flat}V_{g_{0}})$. This set is discrete: for an Anosov flow the set of periods of periodic orbits is a discrete subset of $(0,\infty)$ (see, for instance, [5, Ch. 18]). Hence a continuous map into this set is locally constant, and since $(-\varepsilon,\varepsilon)$ is connected it is constant on the whole interval. Therefore
(26) $L_{[\alpha]}(t)\equiv L_{[\alpha]}(0)\qquad\text{for all free homotopy classes }[\alpha].$
In particular, for every closed geodesic $\gamma$ and every $m\geq 1$,
(27) $\dot{L}_{\gamma^{m}}=0.$
By (17), this is equivalent to the vanishing of the $h$-period integrals:
[…]
5.3. Proof of Theorem 1.2
Proof.
Fix $t_{0}\in(-\varepsilon,\varepsilon)$ and set $g:=g_{t_{0}}$. Let $p_{t}(x,\xi)=|\xi|_{g_{t}}$ and let $X_{g}$ denote the generator of the $g$-geodesic flow on the unit cosphere bundle $S^{*}_{g}X=\{p_{t_{0}}=1\}$. By §5.2, every closed $g_{t}$-geodesic has $t$-independent length; equivalently, for each periodic orbit $y_{t}$ of the Hamiltonian flow of $p_{t}$ on $\{p_{t}=1\}$, its period is independent of $t$.
Let $\dot{p}:=\partial_{t}p_{t}|_{t=t_{0}}$, viewed as a smooth function on $S^{*}_{g}X$. Applying Lemma 5.1 to the shifted family $t\mapsto p_{t_{0}+t}$ shows that for every closed orbit $y_{0}$ of $X_{g}$,
(29) $\int_{y_{0}}\dot{p}\,ds=0.$
Since $g$ has negative curvature, the flow of $X_{g}$ is Anosov. Therefore Livšic’s theorem applies: the vanishing of all periodic orbit integrals (29) implies that $\dot{p}$ is a coboundary, i.e. there exists a (Hölder) function $u$ on $S^{*}_{g}X$ such that
(30) $X_{g}u=\dot{p}$
(for general Anosov flows, smooth regularity of the Livšic coboundary follows from [7] and [4]. In our setting, smoothness follows independently from the $SO(2)$-mode reduction below).
Now, $\dot{p}$ is even under the flip $\xi\mapsto-\xi$ and has $SO(2)$-Fourier support in modes $0,\pm 2$ (cf. (19)). The Guillemin-Kazhdan $SO(2)$-calculus on the coframe bundle (Appendix 6.2, cf. [3]) upgrades the Livšic solution: one can choose $u\in C^{\infty}(S^{*}_{g}X)$ solving (30), and moreover $u$ is fiber-linear (its Fourier support is contained in modes $\pm 1$). Concretely, there is a smooth vector field $v$ on $X$ such that
(31) $u(x,\xi)=\langle\xi,v(x)\rangle_{g}\qquad\text{for }(x,\xi)\in S^{*}_{g}X.$
To convert (30) into a statement about the metric variation, identify $\xi\in S^{*}_{g}X$ with its dual unit tangent vector $T:=\xi^{\sharp}\in S_{g}X$. Then $u=g(v,T)$. Along any $g$-geodesic, $\nabla_{T}T=0$, and therefore
$X_{g}u=\frac{d}{ds}g(v,T)=g(\nabla_{T}v,T)=\frac{1}{2}(\mathcal{L}_{v}g)(T,T).$
On the other hand, (19) at time $t=t_{0}$ reads $\dot{p}=-\frac{1}{2}(\partial_{t}g_{t}|_{t=t_{0}})(T,T)$. Comparing with (30) gives
[…]
hence
(32) $\partial_{t}g_{t}\big|_{t=t_{0}}=-\,\mathcal{L}_{v}g_{t_{0}}.$
Applying the same argument at each $t_{0}$ produces a smooth time-dependent vector field $v_{t}$ on $X$ such that $\partial_{t}g_{t}=-\mathcal{L}_{v_{t}}g_{t}$ for all $t$ in a possibly smaller interval. Let $\phi_{t}$ be the flow of $v_{t}$, i.e. the unique solution of
$\frac{d}{dt}\phi_{t}=v_{t}\circ\phi_{t},\qquad\phi_{0}=\mathrm{id}.$
Then
$\frac{d}{dt}\big(\phi_{t}^{*}g_{t}\big)=\phi_{t}^{*}\big(\partial_{t}g_{t}+\mathcal{L}_{v_{t}}g_{t}\big)=0,$
so $\phi_{t}^{*}g_{t}=g_{0}$ for all $t$. Setting $\varphi_{t}:=\phi_{t}^{-1}$ we obtain $g_{t}=\varphi_{t}^{*}g_{0}$, as claimed. ∎
6. Appendix
6.1. Explicit computation of the lower order terms in (7)
We work at a fixed metric $g$ of negative curvature. Recall the definition of the flat trace (cf. (8) with $t$ fixed):
(33) $\mathrm{Tr}^{\flat}V(\tau)=\int_{S^{*}_{g}X}\delta\!\big(\zeta-G^{\tau}\zeta\big)\,d\mu_{L}(\zeta),\qquad\tau>0,$
where $\delta(\zeta-G^{\tau}\zeta)$ is the Dirac distribution on the fixed point set of $G^{\tau}$ (interpreted via the pullback to the diagonal in the sense of clean intersection).
Fix a primitive closed geodesic $\gamma$ of prime length $L^{\#}_{\gamma}$, and its $m$-iterate $\gamma^{m}$ of length $L_{\gamma^{m}}=mL^{\#}_{\gamma}$. Choose a flow box $U$ around $\gamma$ and a smooth diffeomorphism
(34) $\Psi:\ \mathbb{S}^{1}_{L^{\#}_{\gamma}}\times B_{\rho}(0)\subset\mathbb{S}^{1}_{L^{\#}_{\gamma}}\times\mathbb{R}^{2}\longrightarrow U\subset S^{*}_{g}X,\qquad(s,y)\longmapsto\Psi(s,y),$
such that $\Psi(s,0)\in\gamma$ parametrizes $\gamma$ by arclength $s$ and the geodesic flow has the exact normal form
(35) $G^{\tau}\big(\Psi(s,y)\big)=\Psi\big(s+\tau,\ \Phi_{\tau}(y)\big),$
for a smooth family of local diffeomorphisms $\Phi_{\tau}$ of $B_{\rho}(0)$ with $\Phi_{0}=\mathrm{id}$ and
(36) $D\Phi_{mL^{\#}_{\gamma}}(0)=P_{\gamma}^{\,m}.$
Let $\chi\in C_{c}^{\infty}(U)$ be a cutoff which equals $1$ in a smaller neighborhood of $\gamma$ and set $\chi^{\Psi}:=\chi\circ\Psi$. The localized flat trace is the distribution
(37) $\mathrm{Tr}^{\flat}V_{\gamma}(\tau):=\int_{S^{*}_{g}X}\chi(\zeta)\,\delta\!\big(\zeta-G^{\tau}\zeta\big)\,d\mu_{L}(\zeta),\qquad\mathrm{Tr}^{\flat}V=\sum_{\gamma}\mathrm{Tr}^{\flat}V_{\gamma}+C^{\infty}(\tau).$
Reduction to a $\delta$-calculus on $\mathbb{S}^{1}_{L^{\#}_{\gamma}}\times\mathbb{R}^{2}$. Write the Liouville density in flow box coordinates as
(38) $\Psi^{*}(d\mu_{L})=\rho(s,y)\,ds\,dy,\qquad\rho\in C^{\infty},\ \rho>0.$
Since $d\mu_{L}$ restricts to arclength on the closed orbit $\gamma$ and $s$ is arclength, we may (and do) normalize $\Psi$ so that
(39) $\rho(s,0)\equiv 1\quad\text{for all }s\in\mathbb{S}^{1}_{L^{\#}_{\gamma}}.$
Pair $\mathrm{Tr}^{\flat}V_{\gamma}$ against an arbitrary test function $\varphi\in C_{c}^{\infty}((0,\infty))$. By definition,
$\langle\mathrm{Tr}^{\flat}V_{\gamma},\varphi\rangle$ $=\int_{0}^{\infty}\varphi(\tau)\int_{S^{*}_{g}X}\chi(\zeta)\,\delta\!\big(\zeta-G^{\tau}\zeta\big)\,d\mu_{L}(\zeta)\,d\tau$
(40) $=\int_{0}^{\infty}\!\!\int_{\mathbb{S}^{1}_{L^{\#}_{\gamma}}}\!\!\int_{\mathbb{R}^{2}}\varphi(\tau)\,\chi^{\Psi}(s,y)\,\delta\!\Big(\Psi(s,y)-\Psi(s+\tau,\Phi_{\tau}(y))\Big)\,\rho(s,y)\,dy\,ds\,d\tau.$
Because $\Psi$ is a diffeomorphism, the distribution $\delta(\Psi(s,y)-\Psi(s^{\prime},y^{\prime}))$ is the pullback of the delta on the diagonal under $(\Psi\times\Psi)$, and the Jacobian factors cancel against $\Psi^{*}(d\mu_{L})$ in (40). Consequently, in the variables $(s,y)$ the delta constraint is exactly the simultaneous constraint
(41) $s+\tau\equiv s\ \ (\mathrm{mod}\ L^{\#}_{\gamma}),\qquad\Phi_{\tau}(y)=y,$
and we may rewrite (40) as the $\delta$-pairing for the map $F(s,y,\tau):=(s-(s+\tau),\,y-\Phi_{\tau}(y))$ on $\mathbb{S}^{1}_{L^{\#}_{\gamma}}\times\mathbb{R}^{2}\times\mathbb{R}$:
(42) $\langle\mathrm{Tr}^{\flat}V_{\gamma},\varphi\rangle=\int_{0}^{\infty}\!\!\int_{\mathbb{S}^{1}_{L^{\#}_{\gamma}}}\!\!\int_{\mathbb{R}^{2}}\varphi(\tau)\,\chi^{\Psi}(s,y)\,\rho(s,y)\,\delta_{\mathbb{S}^{1}_{L^{\#}_{\gamma}}}\!\big(s-(s+\tau)\big)\,\delta_{\mathbb{R}^{2}}\!\big(y-\Phi_{\tau}(y)\big)\,dy\,ds\,d\tau.$
The circle delta and the Dirac comb in $\tau$. Let $\delta_{\mathbb{S}^{1}_{L^{\#}_{\gamma}}}$ denote the delta distribution on the circle of length $L^{\#}_{\gamma}$. Its pullback to $\mathbb{R}$ is the Dirac comb
(43) $\delta_{\mathbb{S}^{1}_{L^{\#}_{\gamma}}}(u)=\sum_{k\in\mathbb{Z}}\delta(u-kL^{\#}_{\gamma})\qquad\text{in }\mathcal{D}^{\prime}(\mathbb{R}).$
We have
[…]
Substituting (44) into (42) reduces the pairing to
(45) $\langle\mathrm{Tr}^{\flat}V_{\gamma},\varphi\rangle=\int_{0}^{\infty}\varphi(\tau)\,\Big(L^{\#}_{\gamma}\sum_{k\in\mathbb{Z}}\delta(\tau-kL^{\#}_{\gamma})\Big)\,\Big[\int_{\mathbb{R}^{2}}A(\tau,y)\,\delta\!\big(y-\Phi_{\tau}(y)\big)\,dy\Big]\,d\tau,$
where
(46) $A(\tau,y):=\int_{\mathbb{S}^{1}_{L^{\#}_{\gamma}}}\chi^{\Psi}(s,y)\,\rho(s,y)\,ds.$
Because $\chi^{\Psi}\equiv 1$ near $y=0$ and (39) holds, we have
(47) $A(\tau,0)=\int_{\mathbb{S}^{1}_{L^{\#}_{\gamma}}}1\,ds=L^{\#}_{\gamma}\qquad\text{for $\tau$ in a neighborhood of $mL^{\#}_{\gamma}$.}$
The transversal delta and the Poincaré determinant. Fix $m\geq 1$ and localize in $\tau$ to a small neighborhood of $mL^{\#}_{\gamma}$ so that $\gamma^{m}$ is the unique fixed orbit in the support of $\chi$. Set
(48) $F_{\tau}(y):=y-\Phi_{\tau}(y).$
At $\tau=mL^{\#}_{\gamma}$, $F_{mL^{\#}_{\gamma}}(0)=0$ and, by (36),
(49) $DF_{mL^{\#}_{\gamma}}(0)=I-D\Phi_{mL^{\#}_{\gamma}}(0)=I-P_{\gamma}^{\,m}.$
Negative curvature implies $\gamma$ is hyperbolic, hence $1$ is not an eigenvalue of $P_{\gamma}^{\,m}$ and $\det(I-P_{\gamma}^{\,m})\neq 0$. By the inverse function theorem, $F_{mL^{\#}_{\gamma}}$ is a diffeomorphism from a neighborhood of 0 onto a neighborhood of 0, and the distribution $\delta(F_{mL^{\#}_{\gamma}}(y))$ is computed by a change of variables: for any $\psi\in C_{c}^{\infty}(\mathbb{R}^{2})$ supported sufficiently close to 0,
$\int_{\mathbb{R}^{2}}\psi(y)\,\delta\!\big(F_{mL^{\#}_{\gamma}}(y)\big)\,dy$ $=\int_{\mathbb{R}^{2}}\psi\big(F_{mL^{\#}_{\gamma}}^{-1}(z)\big)\,\delta(z)\,\big|\det DF_{mL^{\#}_{\gamma}}(F_{mL^{\#}_{\gamma}}^{-1}(z))\big|^{-1}\,dz$
(50) $=\psi(0)\,\big|\det DF_{mL^{\#}_{\gamma}}(0)\big|^{-1}=\psi(0)\,\big|\det(I-P_{\gamma}^{\,m})\big|^{-1}.$
Applying (50) to the inner bracket in (45) with $\psi(y)=A(\tau,y)$ and using (47) yields, near $\tau=mL^{\#}_{\gamma}$,
(51) $\int_{\mathbb{R}^{2}}A(\tau,y)\,\delta\!\big(y-\Phi_{mL^{\#}_{\gamma}}(y)\big)\,dy=\frac{A(mL^{\#}_{\gamma},0)}{|\det(I-P_{\gamma}^{\,m})|}=\frac{L^{\#}_{\gamma}}{|\det(I-P_{\gamma}^{\,m})|}.$
Substituting (51) into (45) and restricting to $\tau>0$ gives
[…]
hence, as distributions on $(0,\infty)$,
(53) $\mathrm{Tr}^{\flat}V_{\gamma}(\tau)=\sum_{m\geq 1}\frac{L^{\#}_{\gamma}}{|\det(I-P_{\gamma}^{\,m})|}\,\delta(\tau-mL^{\#}_{\gamma})\quad\text{microlocally near }\tau=mL^{\#}_{\gamma}$.
Conclusion. Summing (53) over all primitive closed geodesics and adding the $C^{\infty}$ contribution from $\tau$ away from the length spectrum yields the sharp Lefschetz flat-trace formula
(54) $\mathrm{Tr}^{\flat}V(\tau)=\sum_{\gamma}\sum_{m\geq 1}\frac{L^{\#}_{\gamma}}{|\det(I-P_{\gamma}^{\,m})|}\,\delta(\tau-L_{\gamma^{m}})\;+\;C^{\infty}(\tau).$
In particular, the parenthetical term “(lower order terms at $\tau=L_{\gamma^{m}}$)” in (7) vanishes identically in the Lefschetz (hyperbolic) setting: there are no additional singular contributions supported at $\tau=L_{\gamma^{m}}$ beyond the Dirac masses already displayed.
6.2. The $SO(2)$ Fourier calculus on $S^{*}_{g}X$ and the $\pm 1$ mode reduction
Let $(X,g)$ be a closed oriented surface with Gaussian curvature $K<0$, and identify $S^{*}_{g}X\simeq SX$ via the musical isomorphism. Denote by $\pi:SX\to X$ the bundle projection. We use the standard canonical coframing on $SX$: there exist smooth $1$-forms $(\alpha,\beta,\psi)$ and smooth vector fields $(X,X_{\perp},V)$ on $SX$ uniquely specified by
$\alpha(X)=1,\ \beta(X)=0,\ \psi(X)=0;\quad\alpha(X_{\perp})=0,\ \beta(X_{\perp})=1,\ \psi(X_{\perp})=0;$
(55) $\alpha(V)=\beta(V)=0,\psi(V)=1,$
together with the structure equations
(56) $d\alpha=\psi\wedge\beta,\qquad d\beta=-\psi\wedge\alpha,\qquad d\psi=K\,\alpha\wedge\beta.$
The Liouville volume form is
(57) $d\mu=\alpha\wedge d\alpha=\alpha\wedge\psi\wedge\beta.$
The geodesic flow generator is $X$ (the Reeb field of $\alpha$), $V$ generates the right $SO(2)$-action (rotation of the fiber angle), and $X_{\perp}=[V,X]$ is the horizontal rotation by $\pi/2$.
Commutators. For any $1$-form $\omega$ and vector fields $Y,Z$,
(58) $d\omega(Y,Z)=Y(\omega(Z))-Z(\omega(Y))-\omega([Y,Z]).$
Applying (58) to $\omega=\alpha,\beta,\psi$ and the pairs $(X,V)$, $(X_{\perp},V)$, $(X,X_{\perp})$, using (55)-(56), gives:
$d\alpha(X,V)$ $=(\psi\wedge\beta)(X,V)=\psi(X)\beta(V)-\psi(V)\beta(X)=0-1\cdot 0=0\ \Rightarrow\ \alpha([X,V])=0,$
$d\beta(X,V)$ $=(-\psi\wedge\alpha)(X,V)=-(\psi(X)\alpha(V)-\psi(V)\alpha(X))=-(0-1\cdot 1)=1\ \Rightarrow\ \beta([X,V])=-1,$
$d\psi(X,V)$ $=(K\alpha\wedge\beta)(X,V)=K(\alpha(X)\beta(V)-\alpha(V)\beta(X))=0\ \Rightarrow\ \psi([X,V])=0,$
hence $X,V]=-X_{\perp}$. Similarly, evaluating at $(X_{\perp},V)$ yields $V,X_{\perp}]=-X$. Finally, at $(X,X_{\perp})$,
$d\psi(X,X_{\perp})=K\alpha\wedge\beta(X,X_{\perp})=K,\qquad d\psi(X,X_{\perp})=-\psi([X,X_{\perp}]),$
so $\psi([X,X_{\perp}])=-K$, and since $\alpha([X,X_{\perp}])=\beta([X,X_{\perp}])=0$ (from $d\alpha(X,X_{\perp})=d\beta(X,X_{\perp})=0$), we obtain
(59) $V,X]=X_{\perp},\qquad[V,X_{\perp}]=-X,\qquad[X,X_{\perp}]=-K\,V.$
Skew-adjointness. We claim that $X,X_{\perp},V$ are divergence-free with respect to $d\mu$, hence skew-adjoint on $L^{2}(SX,d\mu)$. By Cartan’s formula $L_{Y}\omega=i_{Y}d\omega+d(i_{Y}\omega)$ and (56), we have
$L_{X}\alpha=i_{X}(\psi\wedge\beta)+d(\alpha(X))=\psi(X)\beta-\beta(X)\psi+0=0,\qquad L_{X}d\alpha=d(L_{X}\alpha)=0,$
thus $L_{X}(\alpha\wedge d\alpha)=0$. Next,
$L_{X_{\perp}}\alpha=i_{X_{\perp}}(\psi\wedge\beta)=\psi(X_{\perp})\beta-\beta(X_{\perp})\psi=-\psi,\qquad L_{X_{\perp}}d\alpha=d(-\psi)=-d\psi=-K\,\alpha\wedge\beta,$
and therefore
$L_{X_{\perp}}(\alpha\wedge d\alpha)=(-\psi)\wedge d\alpha+\alpha\wedge(-K\alpha\wedge\beta)=0.$
Finally,
$L_{V}\alpha=i_{V}(\psi\wedge\beta)=\psi(V)\beta-\beta(V)\psi=\beta,\qquad L_{V}d\alpha=d\beta=-\psi\wedge\alpha,$
so
$L_{V}(\alpha\wedge d\alpha)=\beta\wedge d\alpha+\alpha\wedge(-\psi\wedge\alpha)=0.$
Hence $L_{X}d\mu=L_{X_{\perp}}d\mu=L_{V}d\mu=0$, and by integration by parts,
(60) $\langle Xu,v\rangle=-\langle u,Xv\rangle,\qquad\langle X_{\perp}u,v\rangle=-\langle u,X_{\perp}v\rangle,\qquad\langle Vu,v\rangle=-\langle u,Vv\rangle,$
for all smooth $u,v$ (and by density for all $H^{1}$ functions).
Fourier decomposition and the operators $\eta^{\pm}$. Let $V$ be the infinitesimal generator of the right $SO(2)$-action; in a fiber angle coordinate $\theta$ one has $V=\partial_{\theta}$, hence $V$ is skew-adjoint and its spectrum is $\{im:\ m\in\mathbb{Z}\}$. Define
$\mathcal{H}_{m}:=\{u\in L^{2}(SX):\ Vu=imu\ \text{in }\mathcal{D}^{\prime}(SX)\},\qquad L^{2}(SX)=\widehat{\bigoplus}_{m\in\mathbb{Z}}\mathcal{H}_{m},\qquad u=\sum_{m\in\mathbb{Z}}u_{m}.$
Introduce the raising/lowering operators
(61) $\eta^{+}:=\tfrac{1}{2}(X-iX_{\perp}),\qquad\eta^{-}:=\tfrac{1}{2}(X+iX_{\perp}),\qquad X=\eta^{+}+\eta^{-}.$
By (60), $(\eta^{+})^{*}=-\eta^{-}$ and $(\eta^{-})^{*}=-\eta^{+}$. Moreover, by (59),
(62) $\eta^{-},\eta^{+}]=-\tfrac{i}{2}[X,X_{\perp}]=-\tfrac{i}{2}(-KV)=\tfrac{i}{2}\,K\,V,\qquad[V,\eta^{\pm}]=\pm i\,\eta^{\pm},$
so $\eta^{\pm}:\mathcal{H}_{m}\to\mathcal{H}_{m\pm 1}$.
Parity reduction for even data. Let $\mathcal{A}:SX\to SX$ be the flip $\mathcal{A}(x,v)=(x,-v)$. Then $\mathcal{A}$ conjugates the flow to its inverse:
(63) $G^{\tau}\circ\mathcal{A}=\mathcal{A}\circ G^{\tau}_{-1},\qquad\tau\in\mathbb{R},$
hence $\mathcal{A}_{*}X=-X$, while $\mathcal{A}$ is a fiber rotation by $\pi$ and therefore commutes with $V$. In particular, if $f$ is even under the flip, $f\circ\mathcal{A}=f$, and $Xu=f$, then
$X(u\circ\mathcal{A})=(\mathcal{A}_{*}X)u\circ\mathcal{A}=-(Xu)\circ\mathcal{A}=-f,$
so $X(u+u\circ\mathcal{A})=0$. Since the geodesic flow on a negatively curved surface is ergodic with respect to $d\mu$, any $L^{2}$ function annihilated by $X$ is constant a.e.; for continuous functions this implies global constancy. Subtracting half this constant from $u$, we may assume
(64) $u\circ\mathcal{A}=-u,$
i.e. $u(x,\theta+\pi)=-u(x,\theta)$ in an angle coordinate. Consequently, all even Fourier modes vanish:
(65) $u_{m}=0\quad\text{for every even }m.$
Mode-by-mode form of the Livšic equation. Let $u\in C^{1}(SX)$ solve
(66) $Xu=f,$
with $f\in C^{\infty}(SX)$. Decomposing $u=\sum u_{m}$ and $f=\sum f_{m}$ and using $X=\eta^{+}+\eta^{-}$ and $\eta^{\pm}:\mathcal{H}_{m}\to\mathcal{H}_{m\pm 1}$ gives, by projection onto $\mathcal{H}_{m}$,
(67) $f_{m}=\eta^{+}u_{m-1}+\eta^{-}u_{m+1},\qquad m\in\mathbb{Z}.$
In our application $f=\dot{p}|_{SX}=-\tfrac{1}{2}h(v,v)$, where $h=\dot{g}$ is a smooth symmetric $2$-tensor on $X$ and $v\in S_{x}X$ is the unit direction. To determine the $SO(2)$-Fourier support of $\dot{p}$, fix a local oriented $g$-orthonormal frame $(e_{1},e_{2})$ on $X$ and write the fiber angle coordinate $\theta$ so that
$v=\cos\theta\,e_{1}+\sin\theta\,e_{2}.$
Set $h_{ij}(x):=h_{x}(e_{i},e_{j})$. Then, using $\cos^{2}\theta=\frac{1}{2}(1+\cos 2\theta)$, $\sin^{2}\theta=\frac{1}{2}(1-\cos 2\theta)$, and $\sin\theta\cos\theta=\frac{1}{2}\sin 2\theta$, we compute
$h(v,v)$ $=h_{11}\cos^{2}\theta+2h_{12}\sin\theta\cos\theta+h_{22}\sin^{2}\theta$
$=\tfrac{1}{2}(h_{11}+h_{22})+\tfrac{1}{2}(h_{11}-h_{22})\cos 2\theta+h_{12}\sin 2\theta.$
Equivalently, writing $\cos 2\theta=\frac{1}{2}(e^{2i\theta}+e^{-2i\theta})$ and $\sin 2\theta=\frac{1}{2i}(e^{2i\theta}-e^{-2i\theta})$, we obtain
(68) $\dot{p}(x,\theta)=-\frac{1}{4}\,(h_{11}+h_{22})(x)-\frac{1}{8}\Big((h_{11}-h_{22})(x)-2i\,h_{12}(x)\Big)e^{2i\theta}-\frac{1}{8}\Big((h_{11}-h_{22})(x)+2i\,h_{12}(x)\Big)e^{-2i\theta}.$
In particular,
$\dot{p}\in\mathcal{H}_{-2}\oplus\mathcal{H}_{0}\oplus\mathcal{H}_{2},$
and $\dot{p}$ is even under the flip $\mathcal{A}(x,v)=(x,-v)$.
Energy-curvature identity. For $w\in C^{\infty}(SX)$, using $(\eta^{\pm})^{*}=-\eta^{\mp}$ and (62),
$\|\eta^{+}w\|_{L^{2}}^{2}-\|\eta^{-}w\|_{L^{2}}^{2}$ $=\langle w,(\eta^{+})^{*}\eta^{+}w\rangle-\langle w,(\eta^{-})^{*}\eta^{-}w\rangle$
$=\langle w,-\eta^{-}\eta^{+}w\rangle-\langle w,-\eta^{+}\eta^{-}w\rangle=\langle w,[\eta^{+},\eta^{-}]w\rangle=-\langle w,[\eta^{-},\eta^{+}]w\rangle$
(69) $=-\Big\langle w,\tfrac{i}{2}KVw\Big\rangle=-\tfrac{i}{2}\int_{SX}K\,(Vw)\,\overline{w}\,d\mu.$
If $w\in\mathcal{H}_{m}$ so that $Vw=imw$, then (69) becomes
(70) $\|\eta^{+}w\|^{2}-\|\eta^{-}w\|^{2}=\frac{m}{2}\int_{SX}K\,|w|^{2}\,d\mu.$
Since $K\leq-\kappa_{0}<0$, for $m>0$ we obtain the coercive estimate
(71) $\|\eta^{+}w\|^{2}\leq\|\eta^{-}w\|^{2}-\frac{\kappa_{0}\,m}{2}\,\|w\|^{2}.$
The $\pm 1$-mode reduction. Assume $u\in C^{1}(SX)$ satisfies $Xu=f$ with $f$ even and supported in modes $0,\pm 2$, and normalize $u$ to be odd as in (64), so $u$ has only odd modes. Then $f_{m}=0$ for all odd $m$ and (67) yields, for all $|m|\geq 3$,
(72) $\eta^{+}u_{m-1}+\eta^{-}u_{m+1}=0.$
Taking $L^{2}$ norms and using orthogonality of Fourier modes implies
(73) $\|\eta^{-}u_{m+1}\|=\|\eta^{+}u_{m-1}\|\qquad(|m|\geq 3),$
and therefore, for $m\geq 3$,
(74) $\|\eta^{+}u_{m}\|^{2}-\|\eta^{-}u_{m}\|^{2}=\|\eta^{-}u_{m+2}\|^{2}-\|\eta^{+}u_{m-2}\|^{2}.$
Summing (74) for $m=3,4,\dots,M$ gives the telescoping identity
(75) $\sum_{m=3}^{M}\big(\|\eta^{+}u_{m}\|^{2}-\|\eta^{-}u_{m}\|^{2}\big)=\|\eta^{-}u_{M+2}\|^{2}+\|\eta^{-}u_{M+1}\|^{2}-\|\eta^{+}u_{1}\|^{2}-\|\eta^{+}u_{2}\|^{2}.$
By (70) and $K\leq-\kappa_{0}$,
(76) $\sum_{m=3}^{M}\big(\|\eta^{+}u_{m}\|^{2}-\|\eta^{-}u_{m}\|^{2}\big)=\sum_{m=3}^{M}\frac{m}{2}\int K|u_{m}|^{2}\,d\mu\leq-\frac{\kappa_{0}}{2}\sum_{m=3}^{M}m\,\|u_{m}\|^{2}.$
Since $\eta^{-}u\in L^{2}(SX)$ and $\eta^{-}u=\sum_{m}\eta^{-}u_{m}$ with orthogonal summands (because $\eta^{-}u_{m}\in\mathcal{H}_{m-1}$), we have $\|\eta^{-}u_{m}\|\to 0$ as $|m|\to\infty$, hence letting $M\to\infty$ in (75)-(76) yields
$\|\eta^{+}u_{1}\|^{2}-\|\eta^{+}u_{2}\|^{2}\leq-\frac{\kappa_{0}}{2}\sum_{m\geq 3}m\,\|u_{m}\|^{2}.$
Both sides are $\leq 0$, so necessarily $\sum_{m\geq 3}m\,\|u_{m}\|^{2}=0$, hence $u_{m}\equiv 0$ for all $m\geq 3$. Applying the same argument to the negative tail (using (72) for $m\leq-3$) gives $u_{m}\equiv 0$ for all $m\leq-3$. Together with (65) we conclude
(77) $u=u_{-1}+u_{1},\qquad u_{\pm 1}\in\mathcal{H}_{\pm 1}.$
Elliptic regularity and fiber-linearity. With (77), the $m=\pm 2$ components of (67) read
(78) $f_{2}=\eta^{+}u_{1},\qquad f_{-2}=\eta^{-}u_{-1}.$
The operators $\eta^{+}:\mathcal{H}_{1}\to\mathcal{H}_{2}$ and $\eta^{-}:\mathcal{H}_{-1}\to\mathcal{H}_{-2}$ are first-order elliptic (their principal symbols are nonvanishing complex combinations of the horizontal symbols of $X$ and $X_{\perp}$), hence by elliptic regularity $f_{\pm 2}\in C^{\infty}$ implies $u_{\pm 1}\in C^{\infty}$. Finally, $\mathcal{H}_{\pm 1}$ consists precisely of fiber-linear functions: in a local oriented orthonormal frame $(e_{1},e_{2})$ and fiber angle $\theta$,
$u_{1}(x,\theta)=a(x)e^{i\theta},\qquad u_{-1}(x,\theta)=\overline{a(x)}e^{-i\theta},$
so $u=u_{-1}+u_{1}=v^{1}(x)\cos\theta+v^{2}(x)\sin\theta$ with a unique smooth vector field $v=v^{1}e_{1}+v^{2}e_{2}$ on $X$. Equivalently,
(79) $u(x,\xi)=\langle\xi,\ v(x)\rangle_{g}\qquad\text{on }S^{*}_{g}X.$
This is the $\pm 1$-mode reduction used in (32).

### v2

@@
 This completes a proof of infinitesimal triviality which does not appeal to marked length spectrum rigidity and which is expected to be robust for Anosov flows in higher dimension.
-Remark 1.1.
+Remark $1.$
 In dimension $n\geq 3$, the Lefschetz fixed point reduction and the Abel-type extraction of the $\delta^{\prime}$-term remain formally identical, but the clean codimension and the harmonic bookkeeping change.
@@
 The rigidity statement in Theorem 1.2 is weaker than marked length spectrum rigidity, since it assumes a smooth deformation.
-The point of the present paper is the distributional first variation formula for the flat trace, in particular the transport coefficient in Proposition 2.2.
-This coefficient does not seem to appear in the literature in this explicit form, and it is the mechanism that converts constancy of a zeta type invariant into infinitesimal constraints on the marked lengths.
+The point of the present paper is the distributional first variation formula for the flat trace, in particular the transport coefficient in Proposition $2.$ This coefficient does not seem to appear in the literature in this explicit form, and it is the mechanism that converts constancy of a zeta type invariant into infinitesimal constraints on the marked lengths.
 The appendices collect auxiliary calculations and are included for completeness.
@@
 In negative curvature, the fixed set of $G^{\tau}_{t}$ is clean precisely when $\tau$ equals the period of a (possibly iterated) closed geodesic.
-2.2. Normal form near a closed orbit
+$2.$ Normal form near a closed orbit
 Fix a primitive closed geodesic $\gamma$ of $(X,g_{0})$ with prime length $L^{\#}_{\gamma}>0$, linearized Poincaré map $P_{\gamma}$ and let $\gamma^{m}$ denote its $m$-iterate of length $L_{\gamma^{m}}=mL^{\#}_{\gamma}$.
@@
 Therefore,
-Proposition 2.2.
+Proposition $2.$
 At each isolated length $\ell>0$, the length-transport coefficient of $\delta^{\prime}(\tau-\ell)$, in $\partial_{t}\big|_{0}\mathrm{Tr}^{\flat}V_{t}(\tau)$ equals
@@
 $\int_{y_{0}}\dot{p}\,ds=0,$
-where $ds$ denotes the Hamiltonian time parameter along $y_{0}$.
+where dsds denotes the Hamiltonian time parameter along $y_{0}$.
 Proof.
-Let $\alpha$ be the canonical $1$-form on $T^{*}X$ and $\omega=d\alpha$ the canonical symplectic form. Let $H_{t}$ be the Hamiltonian vector field of $p_{t}$, characterized by
+Let $\alpha$ be the canonical 11-form on $T^{*}X$ and $\omega=d\alpha$ the canonical symplectic form. Let $H_{t}$ be the Hamiltonian vector field of $p_{t}$, characterized by
 $\iota_{H_{t}}\omega=dp_{t}.$
@@
 $=\int_{\partial([0,\varepsilon]\times(\mathbb{R}/L\mathbb{Z}))}\Phi^{*}\alpha=\int_{y_{\varepsilon}}\alpha-\int_{y_{0}}\alpha.$
-Since $p_{t}$ is homogeneous of degree $1$ in $\xi$, Euler’s identity gives $\alpha(H_{t})=p_{t}$, and therefore
+Since $p_{t}$ is homogeneous of degree 11 in $\xi$, Euler’s identity gives $\alpha(H_{t})=p_{t}$, and therefore
 $\int_{y_{t}}\alpha=\int_{0}^{L}\alpha(H_{t})\,ds=\int_{0}^{L}p_{t}\,ds=L,$
@@
 (36) $D\Phi_{mL^{\#}_{\gamma}}(0)=P_{\gamma}^{\,m}.$
-Let $\chi\in C_{c}^{\infty}(U)$ be a cutoff which equals $1$ in a smaller neighborhood of $\gamma$ and set $\chi^{\Psi}:=\chi\circ\Psi$. The localized flat trace is the distribution
+Let $\chi\in C_{c}^{\infty}(U)$ be a cutoff which equals 11 in a smaller neighborhood of $\gamma$ and set $\chi^{\Psi}:=\chi\circ\Psi$. The localized flat trace is the distribution
 (37) $\mathrm{Tr}^{\flat}V_{\gamma}(\tau):=\int_{S^{*}_{g}X}\chi(\zeta)\,\delta\!\big(\zeta-G^{\tau}\zeta\big)\,d\mu_{L}(\zeta),\qquad\mathrm{Tr}^{\flat}V=\sum_{\gamma}\mathrm{Tr}^{\flat}V_{\gamma}+C^{\infty}(\tau).$
@@
 (49) $DF_{mL^{\#}_{\gamma}}(0)=I-D\Phi_{mL^{\#}_{\gamma}}(0)=I-P_{\gamma}^{\,m}.$
-Negative curvature implies $\gamma$ is hyperbolic, hence $1$ is not an eigenvalue of $P_{\gamma}^{\,m}$ and $\det(I-P_{\gamma}^{\,m})\neq 0$.
+Negative curvature implies $\gamma$ is hyperbolic, hence 11 is not an eigenvalue of $P_{\gamma}^{\,m}$ and $\det(I-P_{\gamma}^{\,m})\neq 0$.
 By the inverse function theorem, $F_{mL^{\#}_{\gamma}}$ is a diffeomorphism from a neighborhood of 0 onto a neighborhood of 0, and the distribution $\delta(F_{mL^{\#}_{\gamma}}(y))$ is computed by a change of variables: for any $\psi\in C_{c}^{\infty}(\mathbb{R}^{2})$ supported sufficiently close to 0,
@@
 Denote by $\pi:SX\to X$ the bundle projection.
-We use the standard canonical coframing on $SX$: there exist smooth $1$-forms $(\alpha,\beta,\psi)$ and smooth vector fields $(X,X_{\perp},V)$ on $SX$ uniquely specified by
+We use the standard canonical coframing on SXSX: there exist smooth 11-forms $(\alpha,\beta,\psi)$ and smooth vector fields $(X,X_{\perp},V)$ on SXSX uniquely specified by
 $\alpha(X)=1,\ \beta(X)=0,\ \psi(X)=0;\quad\alpha(X_{\perp})=0,\ \beta(X_{\perp})=1,\ \psi(X_{\perp})=0;$
@@
 The geodesic flow generator is $X$ (the Reeb field of $\alpha$), $V$ generates the right $SO(2)$-action (rotation of the fiber angle), and $X_{\perp}=[V,X]$ is the horizontal rotation by $\pi/2$.
-Commutators. For any $1$-form $\omega$ and vector fields $Y,Z$,
+Commutators. For any 11-form $\omega$ and vector fields $Y,Z$,
 (58) $d\omega(Y,Z)=Y(\omega(Z))-Z(\omega(Y))-\omega([Y,Z]).$
@@
 (67) $f_{m}=\eta^{+}u_{m-1}+\eta^{-}u_{m+1},\qquad m\in\mathbb{Z}.$
-In our application $f=\dot{p}|_{SX}=-\tfrac{1}{2}h(v,v)$, where $h=\dot{g}$ is a smooth symmetric $2$-tensor on $X$ and $v\in S_{x}X$ is the unit direction.
+In our application $f=\dot{p}|_{SX}=-\tfrac{1}{2}h(v,v)$, where $h=\dot{g}$ is a smooth symmetric 22-tensor on $X$ and $v\in S_{x}X$ is the unit direction.
 To determine the $SO(2)$-Fourier support of $\dot{p}$, fix a local oriented $g$-orthonormal frame $(e_{1},e_{2})$ on $X$ and write the fiber angle coordinate $\theta$ so that